OpenStudy (anonymous):

Find the average rate of change of each function on the interval specified. f(x) = 3x^3 - 1 on (-3,3)

OpenStudy (zzr0ck3r):

what class?

OpenStudy (zzr0ck3r):

|f(-3)-f(3)|/6

OpenStudy (dumbcow):

avg rate of change = slope \[slope = \frac{f(3) -f(-3)}{3-(-3)}\]

OpenStudy (zzr0ck3r):

since this is an even function, and its centered at 0, it will have 0 avg v

OpenStudy (zzr0ck3r):

ahh, ok then just use the formula I showed d=r*t r=d/t distance is given by |f(-3)-f(3)| and t is 6 because that's |3-(-3)|

OpenStudy (jhannybean):

Is it really centered at 0?

OpenStudy (zzr0ck3r):

he said it was x^2 first.

OpenStudy (dan815):